OpenStudy (anonymous):

A ladder leans against a house at a 60° angle to the ground.If the ladder extends to a length of 158 inches, what is the height of the house rounded to the nearest hundredth of an inch?

OpenStudy (anonymous):

take the height of the house as x 158 inches is a hypotenuse sin60=x/158 x=sin60*158 x=136.8 so x=137 inches hope it helps if you need any help with it do ask :)